企業行政演習 (D) 2 単位 (選択)

Internship (D)

【授業目的】この授業は，企業や行政の場において実際的な就業体験を行うことにより，企業等の仕組みや仕事に関する理解を深め，学習意欲を喚起するとともに，高い就業意識の育成を図ることを目的とする．

【授業概要】この授業はある一定期間に会社や行政の場においてインターンシップを実施した時に単位を認定するものである．インターンシップ期間の現地での活動や実施報告書に基づいて評価する．

【授業形式】 [形態] 【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修要件】 [要件] 【履修上の注意】 [注意] 【到達目標】

1. 組織の仕組みや業務の流れ，組織目標を達成するための戦略と実践を理解する．

2. 実社会，職場における人間関係やマナーなどに対する理解を深める．【授業計画】 [計画] 【成績評価基準】取り組み状況，実習レポートおよび実施報告書【教科書】 [教科書] 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216611 【対象学生】開講コース学生のみ履修可能【連絡先】 [連絡先] 【備考】 90 時間分の学修記録 (ポートフォリオ) を提出する．詳細については，指

導教員に相談，指示を仰ぐこと．

Internship (D) 2 units (selection)

Target⟩ This program is provided to increase knowledge and experience in the actual industrial technology in various fields. Through the practical experience, the program will allow students to become discerning and well-balanced researchers or engineers with deeper understanding of technological societies.

Outline⟩ Credits are acknowledged to students who experienced the internship in companies and/or goverment offices. Their fulfilment is evaluated by activity reports during internship.

Style⟩ [形態] �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ [要件] Notice⟩ [注意]

Goal⟩

1. Grasping internal systems of a company, management, action and strategy to achieve the company’s goal. (2)

2. Understanding human relationship and communicative manner in a working

society and a company. Schedule⟩ [計画] Evaluation Criteria⟩ Achievement is evaluated by how you tackled, activity

reports, and an evaluation record submitted by supervisers in company and/or government.

Textbook⟩ [教科書] Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216611

Student⟩ Able to be taken by only specified class(es) Contact⟩ [連絡先] Note⟩ A possible student should submit portfolios which are evidenced for his/her

90 hours’ extensive work. Students are advised to ask details about this lecture to their supervisers.